OpenStudy (watchmath):

Stolen problem from Dinda :D Without calculator, find 8^(-x)+8^x if you are given that 4^x+4^(-x)=7

OpenStudy (anonymous):

2^(2x)+2^(-2x)=7 (2^x+2^-x)^2=9 2^x+2^-x=3; let a=2^x a+1/a=3 (a+1/a)^3=27 a^3+1/a^3+3(a+1/a)=27 a^3+1/a^3=18 but a^3=8^x. so ans is 18

OpenStudy (watchmath):

Awesome! :D

OpenStudy (anonymous):

if u have any more prob. u can post it....i'll try

OpenStudy (watchmath):

Ok, I'll repost my old problem that has no answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have answered your next problem...